OpenStudy (anonymous):

the length of a rugby league field is 52 meters longer than its width "w" . The area of a rugb league field is 8160 square meters. Find the dimensions of the field.

OpenStudy (anonymous):

area = lw l = 52 +w so 8160 = 52 + w

OpenStudy (anonymous):

so subtract 8160 - 52 = w and then use that number as w

OpenStudy (anonymous):

its actually 8160 = (52+w)w = 52w + w^2